# Towards Building Multilingual Language Model for Medicine

### v2

@@ 2.2.2 Machine Rating on Rationale
Figure 4: Comparative analysis on model ratings. In Figure a, BLEU score rating denotes the rating score calculated based on ranking by BLEU score. Human rating refers to rankings provided by humans, while GPT-4 rating refers to rankings generated by GPT-4. Figure b presents the correlation between human rating results and different automatic metrics. $\tau$ is the Kendall rank correlation coefficient while $k$ is the slope of fitted line.
2.3 Ablation Studies on Data Composition
@@
In this section, we present an analysis on the effects of data construction process, as depicted in Table 3. Our ablation studies are carried out both on MMedLM and MMedLM 2 under the full fine-tuning setting, leveraging the InternLM and InternLM 2 framework respectively. Overall, the results observed on the two models are mostly consistent, in the following, we will thus focus our discussion on the results on MMedLM 2.
@@
Here, we distinguish HQ-Data (High-Quality Data) and US-Data (Unspecified Source Data). HQ-Data includes content sourced from books and websites, which has undergone thorough human verification, whereas US-Data is derived from filtering medical-related content from a general corpus. The outcomes, detailed in Table 3, reveal that equipping the model with comprehensive rationales results in an average multiple-choice accuracy increase of 2.42 points, elevating from 56.17 to 58.59. However, further auto-regressive training exclusively on the English segment of MMedC does not yield an overall accuracy improvement. This is primarily because the model overfits the English domain, leading to superior performance in English but shortcomings in other languages (check Supplementary Material C for more details). If expanding the auto-regressive training into the whole multilingual medical corpus, the problem will be relieved, significantly improving the final results. This includes not only boosting the choice accuracy to 51.33 but also enhancing reasoning capabilities by 2.34 and 1.54 points respectively on BLEU-1 and ROUGE-1 scores. Moreover, the inclusion of an automatically gathered US-Data facilitates an additional accuracy boost from 51.33 to 55.01, representing a significant increase of 3.68 points. Gain can also be observed in rationale ability, i.e., 0.59 in BLEU-1 and 0.47 in ROUGE-1.
@@
MMedLM $\mathrm{2}_{en}$ 58.28 46.46 42.99
MMedLM $\mathrm{2}_{ocr}$ 64.43 47.95 44.57
@@ 3 Discussion
3.1 Auto-regressive Training on MMedC is Effective.
@@
As revealed in Table 1, both MMedLM and MMedLM 2 demonstrated significant improvement over their original baseline models, namely, InternLM and InternLM 2, underscoring the effectiveness of MMedC for providing targeted domain-specific enhancements across a wide range of existing open-source language models. In addition, the observed performance boosts serve as an indication that the pre-training corpora of existing LLMs exhibit limitations when faced with multilingual medical contexts. Our findings reinforce the necessity of specialized corpora like MMedC to bridge these gaps.
3.2 Incorporating More Data is Generally Effective.
@@
While exploring how varying data sources influence the outcomes of language model performance, our findings, presented in Table 3, reveal that, the inclusion of high-quality multilingual data (HQ-Data) can lead to significant performance improvements. Additionally, we observe that incorporating data filtered from general contexts, despite its relatively lower quality compared to more explicitly medical-related sources, is also effective. This improvement underscores the value of integrating diverse data types within MMedC, to enrich the model’s understanding of medical information across different languages.
3.3 Incorporating Rationale For Fine-tuning is Effective.
@@
While fine-tuning on MMedBench (trainset), we observed that, the integration of rationale data with multiple-choice prediction, can enhance performance on specific tasks. As shown by Table 3, combining correct answers with their rationales during the supervised fine-tuning phase not only enables LLMs to output rationale sentence, but also results in a noteworthy multiple choice accuracy improvement of 2.33% for InternLM and 2.42% for InternLM 2 on the MMedBench (testset). This indicates that the two tasks are strongly correlated and reinforces the significance training on multi-choice prediction and rationale tasks jointly.
3.4 Strong Foundational LLMs Improve the Final Results.
@@
On MMedBench, we also notice that stronger LLM backbones (commonly released later) generally improve the final results on multilingual medical QA. With the release of more advanced LLMs, their pre-training corpus have been expanded significantly, gradually encompassing more languages. Even though non-English languages constituted a small fraction of the total, the sheer volume of the overall corpus allowed the models to encounter a vast array of multilingual texts during training, significantly enhancing their multilingual capabilities, as seen with the comparison between LLaMA 2 and Mistral, where the later Mistral performs much better. Such enhancement in general multilingual language abilities can also improve the performance after adaptation in medical domain (MMedLM vs. MMedLM 2). This observation encourages us to focus more on building up an open-source dataset for medicine, allowing future works to better leverage the rapid improvement on general LLMs.
3.5 The Impact of Multilingual Medical LLMs.
@@
In this work, we initialize the development of medical multilingual LLMs, which has great research and clinical impacts. In research domain, it can promote the following critical directions:
@@
  • General Medical Artificial Intelligence (GMAI). GMAI [17] commits to developing a multimodal AI model that can be directly applied to a wide range of healthcare scenarios, where LLMs are often used as an human-machine interface [18, 30, 27]. Replacing the English-central LLM with a multilingual one enables to make good use of world-wide data source, thus expanding the available multimodal training data, and improving the representation quality for other modalities.
@@
  • Retrieval Augmented Generation. Hallucination is considered as a major problem with currently existing LLMs, especially for medical domain. One potential solution is to develop retrieval-augmented architectures [32, 33, 15]. The key motivation is that by retrieving and inputting facts from extra knowledge bases into LLMs, the generation output can avoid most fatal fact error. However, until now, most efforts are made in English, greatly limiting retrieval augmentation methods to leverage medical knowledge in other languages. Developing multilingual LLMs can benefit the retrieval process, greatly enriching the potential available knowledge base.
@@
On clinical practice, open-source multilingual medical LLMs can also meet the following demands:
@@
  • Language Barriers. In many healthcare systems, language barriers between patients and healthcare providers can hinder effective communication, leading to misunderstandings, misdiagnoses, and inadequate care, causing high-quality medical resources inaccessible for most people. Multilingual medical LLMs can facilitate real-time translation and interpretation, ensuring that patients can effectively communicate their symptoms and understand their diagnoses and treatment options.
@@
  • Cultural and Law Sensitivity. Multilingual medical LLMs can also be trained to recognize and address cultural or law nuances and sensitivities for different countries in healthcare interactions. Understanding cultural backgrounds and law differences can significantly enhance the trust to medical LLMs, leading to better health outcomes.
@@
  • Education. These models can also be customized for education, especially in regions where there is a shortage of medical educators or resources. By providing educational materials and simulations in multiple languages, medical multilingual LLMs can help standardize medical training and ensure consistent quality of care worldwide.
3.6 Limitations and Future Work.
While our work primarily focuses on constructing a multilingual medical corpus and enhancing the capabilities of LLMs in the medical domain across various languages, we encountered certain limitations. First, there are only six languages in our dataset. In the future, we anticipate expanding to include more languages, such as German and Arabic. Second, considering computational costs, our final model is in 7B scale, we might switch the training progress to a larger architecture that may achieve better results according to the scaling law [11]. Lastly, our final model shares the hallucination flaw with other LLMs. In the future, we aim to investigate more retrieval augmentation methods [32].
@@ 4 Methodology
In this part, we will introduce our methodology in detail. Specifically, in Sec. 4.1, we introduce the construction pipeline for MMedC. In Sect. 4.2, we describe the auto-regressive training procedure. In Sect. 4.3, we discuss the new multilingual medical benchmark, MMedBench, including its curation procedure, evaluation settings and metrics.
@@ 4.1 Large-scale Multilingual Medical Corpus
4.1.1 Filtering Medical-related Content
In the broader landscape of natural language processing, the general NLP community has amassed an extensive array of corpora, such as CommonCrawl, which captures billions of web pages monthly and has been operational for years. Despite medical-related content only constitutes a small fraction of this colossal dataset, its sheer volume presents a valuable opportunity for creating a large-scale, medical-specific corpus with the application of sophisticated auto-filtering techniques.
@@ 4.1.1 Filtering Medical-related Content
if $Lang$ = "Space Delimited" thennormal-$\triangleright$ Split text into words first for space delimited languages
@@
Initialize total keyword length $totalLen\leftarrow 0$
@@
Medical Keyword Count. Medical keyword count quantifies the number of unique medical keywords in the texts. Let $K$ be the set of a priori keywords representing medical terms of interest, and let $T$ denote the entire text corpus under analysis. The set of unique keywords appearing in the text can be formulated as $K_{U}=\{k\mid k\in T\wedge k\in K\}$. The Medical Keyword Count (MKC) is then defined as $\text{MKC}=|K_{U}|$, where ||$\cdot$|| ⋅ | denotes the cardinality of the set.
Keyword Density. Keyword density measures the proportion of text occupied by medical keywords relative to the total text length. This metric is instrumental in identifying texts that, despite their length, only incidentally include medical terms. Let $len(T)$ denote the total number of characters in the text $T$, and $occ(t,T)$ denote the occurrence times of word $t$ in $T$. The keyword density, denoted as $D$, can be formulated as:
@@
With the two metrics, we simply set a threshold bar to filter each sentence. To control the filtering quality, we randomly sampled $100$ sentences per language, and on average, 98 sentences are manually checked as medical-related. The final threshold and filtering ratios are detailed in Supplementary Material B.
4.1.2 Medical Textbooks
@@
In addition to filtering the general language corpus, we also collect dozens of medical textbooks, that represent a rich repository of extensive medical knowledge, underscored by a rigorous publication process that ensures content quality. We have curated a collection exceeding 20,000 books, in line with the methodology outlined in PMC-LLaMA [29]. To extract texts from the books, we adopted Optical Character Recognition (OCR) models, specifically, we used the PaddleOCR tool for its proficiency in handling multiple languages. The OCR process generates a list detailing the coordinates and content of each text box, which is then organized in a left-to-right and top-to-bottom order. Furthermore, to ensure a focus on medical content, we excluded non-essential pages, such as covers, tables of contents, and epilogues, identifying them by their page numbers for removal. Quantitatively, we finally collect 4B tokens for English, 1.1B tokens for Chinese, 0.4B tokens for Russian and 0.3B tokens for French.
4.1.3 Medical Websites
@@
Considering that filtering-based data are based on CommonCrawl, which is randomly scratched and untraceable, to avoid missing some important medical knowledge websites, we further crawled a number medical-related websites as compensatory. We focus on three types of websites. Firstly, we target medical encyclopedias, which offer detailed information on diseases and drugs. While this data is of exceptional quality, it is often limited in quantity and subject to stringent access controls. Secondly, we source content from medical consultation platforms and popular science articles about medicine. These sources, though less technical, provide a wealth of knowledge on medical common sense. Lastly, we expand our data collection to include medical news websites, which allow us to gather a larger volume of unrestricted data and incorporate timely information into our model. This strategy enhanced the model’s ability to understand and respond to current medical events and trends. Collecting data from these varied websites, we compile a comprehensive and diverse medical corpus, encompassing in-depth professional medical knowledge as well as a broad spectrum of general medical information and up-to-date industry insights. The detailed websites are listed and as a result, we get 0.1B tokens for Japanese, 0.05B tokens for Spanish, and 0.1M tokens for French.
4.1.4 Existing Small-scale Multilingual Medical Corpus
@@
Apart from the above newly collected data, we also leverage many existing open-source corpus. Specifically, we utilized the following three datasets: Wikipedia [5], Baidu Baike [3], and UFAL Medical Corpus [8]. For Wikipedia and Baidu Baike, we employ the same filtering methodology outlined in Section 4.1.1 to extract the medical domain corpus, while for UFAL, a medical corpus designed for translation tasks, we use it as it is.

### v3

@@ 2 Results
Figure 4: Comparative analysis on model ratings. a Score bars represent ranked scores under different metrics. BLEU score rating denotes the rating score calculated based on ranking by BLEU score. Human rating refers to rankings provided by humans, while GPT-4 rating refers to rankings generated by GPT-4. b The fitted lines present the correlation between human rating results and different automatic metrics. $\tau$ is the Kendall rank correlation coefficient while $k$ is the slope of fitted line.
Evaluation on Public English Benchmarks
Here, we incorporate additional English instructions (from PMC-LLaMA [48]) into MMed-Llama 3 finetuning, and present comparisons with other existing LLMs on English-only benchmarks. Specifically, there are four widely used multiple-choice question-answering benchmarks, namely, MedQA, MedMCQA, PubMedQA and MMLU (Massive Multitask Language Understanding)-Medical [2, 4, 31]. The details on these benchmarks can be found in Section 4.4. Roughly speaking, MedQA and MedMCQA are clinical exams, mainly assessing the diagnosis or treatment ability, PubMedQA focusing on biomedical academical question-answering and MMLU-Medical is a medical sub-split of MMLU, targeting at assessing the basic knowledge for different medical concepts.
As shown in the Table 3, MMed-Llama 3 demonstrate state-of-the-art performance on English benchmarks, specifically, we obtain $4.5\%$, $4.3\%$ and $2.2\%$ performance gain, respectively on MedQA, MedMCQA and PubMedQA. Similarly, on MMLU, our model can achieve the best performance on most results among the open-source LLMs, even surpasing the strong GPT-3.5 significantly, e.g., 72.59 vs. 67.69.
[b]
Table 3: Multiple-choice accuracy evaluation on various English multiple-choice question-answering benchmarks. We report each model’s accuracy across various tasks separately, with “Avg.” denoting the mean score over nine tasks. Note that, for fairness, all the scores are based on basic generation settings without extra ensembling or prompting strategies, like Chain-of-Thought [20] or self-consistency [46]. Since all the English benchmarks share an official test set, we directly use the scores reported in their original papers for other models. For MedAlpaca, GPT-4, GPT-3.5 and Llama 3, ther scores are based on Open Medical-LLM Leaderboard [34].
Method Size Year MedQA MedMCQA PubMedQA MMLU Avg.
CK MG An PM CB CM
Close-source Models
GPT-3.5 - 2022.11 57.7 72.7 53.8 74.7 74.0 65.9 72.8 72.9 64.7 67.69
GPT-4 - 2023.03 85.8 72.3 70.0 90.2 94.0 84.4 94.5 93.8 83.2 85.36
Flan-PaLM 540B 2022.12 67.6 57.6 79.0 80.4 75.0 63.7 83.8 88.9 76.3 74.70
MedPaLM 2 - 2023.05 86.5 72.3 81.8 88.7 92.0 84.4 95.2 95.8 83.2 86.66
Open-source Models
MedAlpaca 7B 2023.3 41.7 37.5 72.8 57.4 69.0 57.0 67.3 65.3 54.3 58.03
PMC-LLaMA 13B 2023.9 56.4 56.0 77.9 - - - - - - -
MEDITRON 7B 2023.11 57.2 59.2 74.4 64.6 59.9 49.3 55.4 53.8 44.8 57.62
Mistral 7B 2023.12 50.8 48.2 75.4 68.7 71.0 55.6 68.4 68.1 59.5 62.97
Gemma 7B 2024.2 47.2 49.0 76.2 69.8 70.0 59.3 66.2 79.9 60.1 64.19
BioMistral 7B 2024.2 50.6 48.1 77.5 59.9 64.0 56.5 60.4 59.0 54.7 58.97
Llama 3 8B 2024.4 60.9 50.7 73.0 72.1 76.0 63.0 77.2 79.9 64.2 68.56
MMed-Llama 3 (Ours) 8B - 65.4 63.5 80.1 71.3 85.0 69.6 77.6 74.3 66.5 72.59
Ablation Studies on Data Composition
@@
We present an analysis on the effects of dataset construction process, as depicted in Table 4. Our ablation studies are carried out on MMedLM, MMedLM 2 and MMed-Llama 3 under the full fine-tuning setting, leveraging the InternLM, InternLM 2 and Llama 3 as base models. Overall, the results observed on the three models are mostly consistent, in the following, we will thus focus our discussion on MMed-Llama 3.
@@
Here, we distinguish HQ-Data (High-Quality Data) and US-Data (Unspecified Source Data). HQ-Data includes content sourced from books and websites, which has undergone thorough human verification, whereas US-Data is derived from filtering medical-related content from a general corpus. The outcomes, detailed in Table 4, reveal that equipping the model with comprehensive rationales results in an average multiple-choice accuracy increase of 4.06 points, elevating from 58.72 to 62.79. However, further auto-regressive training exclusively on the English segment of MMedC does not yield an overall accuracy improvement. We conjecture this is due to overfitting on English, leading to superior performance in English, but inferior results on other languages (check Supplementary Material E for more details). While expanding the auto-regressive training into the entire multilingual medical corpus, the problem can be largely alleviated, significantly improving the final results. This includes not only boosting the choice accuracy to 64.40, but also enhancing reasoning capabilities by 0.48 and 0.54 points respectively on BLEU-1 and ROUGE-1. Moreover, the inclusion of an automatically gathered US-Data facilitates an additional accuracy boost from 64.40 to 67.75, representing a significant increase of 3.35 points. Performance gains can also be observed in rationale ability, i.e., 0.29 in BLEU-1 and 0.16 in ROUGE-1.
[b]
@@
MMedLM $\mathrm{2}_{\text{en}}$ 58.28 46.46 42.99
MMedLM $\mathrm{2}_{\text{ocr}}$ 64.43 47.95 44.57
@@
MMed-Llama 3
Baseline (Llama 3) 58.72 / /
Baseline (Llama 3) 62.79 46.76 42.84
[…]
MMed-Llama 3 67.75 47.22 43.29
@@ 3 Discussion
In this section, we will first highlight the main empirical conclusions from our experimental results, followed by the potential impact of this work, and finally the existing limitations.
Experimental Results. From our experimental results we can draw the following critical conclusions.
@@
First, auto-regressive training on MMedC is effective. As revealed in Table 1, All MMedLM, MMedLM 2 and MMed-Llama 3 demonstrated significant improvement over their original baseline models, namely, InternLM, InternLM 2 and Llama 3, underscoring the effectiveness of MMedC for providing targeted domain-specific knowledge. In addition, the observed performance boosts serve as an indication that the pre-training corpora of existing LLMs exhibit limitations when faced with multilingual medical contexts. Our findings reinforce the necessity of specialized corpora like MMedC to bridge these gaps.
@@
Second, incorporating More Data is Generally Effective. While exploring how varying data sources affect the outcomes of language model performance, our findings, presented in Table 4, reveal that the inclusion of high-quality multilingual data (HQ-Data) can lead to significant performance improvements. Additionally, we observe that incorporating data filtered from general language corpus, despite its relatively lower quality compared to more explicitly medical-related sources, is also effective. This improvement underscores the value of integrating diverse data types within MMedC.
@@
Third, incorporating rationale for fine-tuning is effective. While fine-tuning on MMedBench (trainset), we observed that, the integration of rationale data with multiple-choice prediction, can enhance performance on specific tasks. As shown by Table 4, combining correct answers with their rationales during the supervised fine-tuning phase not only enables LLMs to output rationale sentence, but also results in a noteworthy multiple choice accuracy improvement of 2.33% for InternLM, 2.42% for InternLM 2 and 4.07% for Llama 3 on the MMedBench (testset). This indicates that the two tasks are strongly correlated and reinforces the significance of training on multi-choice prediction and rationale tasks jointly.
@@
Fourth, strong foundational LLMs improve the final results. On MMedBench, we also notice that stronger LLM backbones (commonly released later) generally improve the final results on multilingual medical QA. With the release of more advanced LLMs, their pre-training corpus has been expanded significantly, gradually encompassing more languages. Even though non-English languages constituted a small fraction of the total, the sheer volume of the overall corpus allowed the models to encounter a vast array of multilingual texts during training, significantly enhancing their multilingual capabilities, as seen with the comparison between Llama 2, Mistral and Llama 3, where the later models all performs much better than the former one. Such enhancement in general multilingual language abilities can also improve the performance after adaptation in medical domain (MMedLM vs. MMedLM 2 vs. MMed-Llama 3). This observation shows we should focus more on building up open-source datasets for medicine, that allows future works to better leverage the rapid improvement of general LLMs.
@@
Research Impacts. Moreover, by initiating the development of multilingual medical LLMs, our work can promote the following critical research directions:
@@
Promote General Medical Artificial Intelligence (GMAI) development. GMAI [28] commits to developing a multimodal AI model that can be directly applied to a wide range of healthcare scenarios, where LLMs are often used as a human-machine interface [29, 49, 44]. Replacing the English-centric LLM with a multilingual one enables to make good use of worldwide data source, thus expanding the available multimodal training data, and improving the representation quality for other modalities.
@@
Improve retrieval augmented generation. Hallucination is considered as a major problem with existing LLMs, especially for medical domain. One potential solution is to develop retrieval-augmented architectures [50, 51, 24]. The key motivation is that by retrieving facts from extra knowledge base, the generated outputs from LLMs can avoid most fatal fact error. However, until now, most efforts have been made in English, greatly limiting retrieval-augmented methods to leverage medical knowledge in other languages. Developing multilingual LLMs can benefit the retrieval process, greatly enriching the potential available knowledge base.
@@
Clinical Impacts. Beyond research impacts, on clinical practice, open-source multilingual medical LLMs can also meet the following demands.
@@
Ease the language barrier. In many healthcare systems, language barriers between patients and healthcare providers can hinder effective communication, leading to misunderstandings, misdiagnoses, and inadequate care, causing high-quality medical resources inaccessible for most people. Multilingual medical LLMs can facilitate real-time translation and interpretation, ensuring that patients can effectively communicate their symptoms and understand their diagnoses and treatment options.
@@
Reduce the cultural and law sensitivity. Multilingual medical LLMs can also be trained to recognize and address cultural or law nuances and sensitivities for different countries in healthcare interactions. Understanding cultural backgrounds and law differences can significantly enhance the trust to medical LLMs, leading to better health outcomes.
@@
Help medical education. These models can also be customized for education, especially in regions where there is a shortage of medical educators or resources. By providing educational materials and simulations in multiple languages, medical multilingual LLMs can help standardize medical training and ensure consistent quality of care worldwide.
Potential Limitations. While our work primarily focuses on constructing a multilingual medical corpus and enhancing the capabilities of LLMs for medicine across various languages, we encountered certain limitations.
First, given that a significant portion of our data is acquired through web crawling, it is inevitable that the corpus may contain inherent biases against certain underprivileged populations. This is a critical challenge in the development of medical Language Models (LLMs) as highlighted in previous research [32]. In the future, we will explore more stringent and comprehensive safety controls on potential bias.
Second, on explainability, although we strive to enhance the model with extra rationale capabilities, to help users to understand the final decisions. It remains under-explored on developing explainability for LLM architectures, like those utilized for convolutional blocks or MLPs [18].
Third, the languages in this dataset do not cover all the world population. In the future, we anticipate expanding to include more languages, for example, German and Arabic. Specifically, the common crawl datasets [8] comprise over 167 languages, with our filtering pipeline, we can efficiently extract medical-related terms by defining specific filtering seed words. In addition, in numerous languages, medical literature is available to support local medical education, and integrating these resources into our approach can further enrich the training corpus. Moreover, as general LLMs become increasingly robust, although they may not accurately answer medical questions across various languages, they can effectively rewrite reference sentences into alternative formats or translate them into other languages, tasks that are comparatively simpler. This capability can serve as an augmentation strategy to enhance data for extremely low-resource languages.
Lastly, considering the computational costs, our final model is in 8B scale, in the future, we will switch the training progress to a larger architecture with retrieval augmentation, which can potentially achieve better results, while alleviating hallucination issues.
@@ 4 Methodology
In this part, we provide details on our methodology. Specifically, in Section 4.1, we introduce the construction pipeline for MMedC. In Section 4.2, we describe the auto-regressive training procedure. In Section 4.3, we discuss the new multilingual medical benchmark, MMedBench, including its curation procedure, evaluation settings and metrics.
@@ 4.1 Large-scale Multilingual Medical Corpus
Filtering Medical-related Content. The first way to obtain medical-related content is using heuristic algorithms for filtering. In the broader landscape of natural language processing, the general NLP community has amassed an extensive array of corpora, such as CommonCrawl, which captures billions of web pages monthly and has been operational for years. Despite that medical-related content only constitutes a small fraction of this colossal dataset, its sheer volume presents a valuable opportunity for creating a large-scale, medical-specific corpus with the application of sophisticated auto-filtering techniques.
@@
if $Lang$ = "Space Delimited" $\triangleright$ Split text into words first for space delimited languages
@@
Initialize total keyword length $L\leftarrow 0$
@@
Medical keyword count quantifies the number of unique medical keywords in the texts. Let $K$ be the set of a priori keywords representing medical terms of interest, and let $T$ denote the entire text corpus under analysis. The set of unique keywords appearing in the text can be formulated as $K_{U}=\{k\mid k\in T\wedge k\in K\}$. The Medical Keyword Count (MKC) is then defined as $\text{MKC}=|K_{U}|$, where ||$\cdot$|| ⋅ | denotes the cardinality of the set.
Keyword density measures the proportion of text occupied by medical keywords relative to the total text length. This metric is instrumental in identifying texts that, despite their length, only incidentally include medical terms. Let $len(T)$ denote the total number of characters in the text $T$, and $occ(t,T)$ denote the occurrence times of word $t$ in $T$. The keyword density, denoted as $D$, can be formulated as:
@@
With the two metrics, we simply set a threshold bar to filter each sentence. To control the filtering quality, we randomly sampled $100$ sentences per language, and on average, 98 sentences are manually checked as medical-related. The final threshold and filtering ratios are detailed in Supplementary Material C.
@@
Medical Textbooks. In addition to filtering the general language corpus, we also collect dozens of medical textbooks, that represent a rich repository of extensive medical knowledge, underscored by a rigorous publication process that ensures content quality. We have curated a collection exceeding 20,000 books, in line with the methodology outlined in PMC-LLaMA [48]. To extract texts from the books, we adopted Optical Character Recognition (OCR) models, specifically, we used the PaddleOCR tool for its proficiency in handling multiple languages. The OCR process generates a list detailing the coordinates and content of each text box, which is then organized in a left-to-right and top-to-bottom order. Furthermore, to ensure a focus on medical content, we excluded non-essential pages, such as covers, tables of contents, and epilogues, identifying them by their page numbers for removal. Quantitatively, we finally collect 4B tokens for English, 1.1B tokens for Chinese, 0.4B tokens for Russian and 0.3B tokens for French.
@@
Medical Websites. Considering that filtering-based data are based on CommonCrawl, which is randomly scratched and untraceable, to avoid missing some important medical knowledge websites, we further crawled a number medical-related websites as compensatory. We focus on three types of websites. Firstly, we target medical encyclopedias, which offer detailed information on diseases and drugs. While this data is of exceptional quality, it is often limited in quantity and subject to stringent access controls. Secondly, we source content from medical consultation platforms and popular science articles about medicine. These sources, though less technical, provide a wealth of knowledge on medical common sense. Lastly, we expand our data collection to include medical news websites, which allow us to gather a larger volume of unrestricted data and incorporate timely information into our model. This strategy enhanced the model’s ability to understand and respond to current medical events and trends. Collecting data from these varied websites, we compile a comprehensive and diverse medical corpus, encompassing in-depth professional medical knowledge as well as a broad spectrum of general medical information and up-to-date industry insights. The detailed websites are listed and as a result, we get 0.1B tokens for Japanese, 0.05B tokens for Spanish, and 0.1M tokens for French.
@@
Existing Small-scale Multilingual Medical Corpus. Apart from the above newly collected data, we also leverage many existing open-source corpus. Specifically, we utilized the following three datasets: Wikipedia [10], Baidu Baike [3], and UFAL Medical Corpus [14]. For Wikipedia and Baidu Baike, we employ the same filtering methodology mentioned before to extract the medical domain corpus, while for UFAL, a medical corpus designed for translation tasks, we use it directly.